OpenStudy (ilovebmth1234):

Any Treca Digital Academy (TDA) Students knowing how to do the lesson Solving Problems with Right triangles can you please help me? :3

OpenStudy (dayakar):

can u write any problem?

OpenStudy (ilovebmth1234):

Jeremy finds a set of plans to build a model airplane. Most, but not all, of the lengths given in the plans are direct measurements. According to the plans, each horizontal stabilizer of the tail wing is made from a right triangle. The length of the leg of the triangle that touches the plane is AB = 8 inches, and cosA=817. To find the number of inches each horizontal stabilizer extends from the plane, find BC.

OpenStudy (ilovebmth1234):

it would be shorter but im in the tutorial and got stuck :/

OpenStudy (james1107):

soory dont know this kind of stuff

OpenStudy (james1107):

@alivejeremy

OpenStudy (ilovebmth1234):

me either lol i got stuck

OpenStudy (alivejeremy):

Give me a min i'm helping someone right now <3

OpenStudy (ilovebmth1234):

alright

OpenStudy (dayakar):



OpenStudy (dayakar):

cos A = AB/AC 817 = AB/8 [ GIVEN VALUES]

OpenStudy (dayakar):

CAN U find AB =?

OpenStudy (ilovebmth1234):

where did you get 817?

OpenStudy (ilovebmth1234):

ohhh its supposed to say 8/17

OpenStudy (dayakar):

cos A = 817 is given in the problem

OpenStudy (ilovebmth1234):

right but it was supposed to say 8/17

OpenStudy (dayakar):

now it becomes 8/17 = AB/8

OpenStudy (ilovebmth1234):

ab/8 would be 1

OpenStudy (dayakar):

forget all this, i ' ll tell u one simple method

OpenStudy (ilovebmth1234):

uhh okay

OpenStudy (dayakar):

 IS IT CLEAR

OpenStudy (dayakar):

do u understand it?

OpenStudy (dayakar):

if u get it i will continue

OpenStudy (ilovebmth1234):

yeah my thing says this is how the tutorial solved it \[\cos A=\frac{ AB }{ AC }=\frac{ 8 }{ 17}\] \[AC=17\] now apply the Pythagorean theorem to find BC \[BC=\sqrt{AC^2-AB^2}\] \[BC=\sqrt{17^2-8^2=}\sqrt{225=15}\]

OpenStudy (ilovebmth1234):

so the answer would be 15

OpenStudy (dayakar):

exactly, u got it

OpenStudy (ilovebmth1234):

right but i just dont understand how they solved it lol